OpenStudy (anonymous):

How do I add fractions with a variable and without a variable on the other one. For example in below

OpenStudy (anonymous):

\[\frac{ 7c+2d }{ 8c } - \frac{ c+2d }{ 2c } + 3\]

OpenStudy (ddcamp):

When adding fractions, you find a common denominator. What could you multiply the top and bottom of the second fraction by so that it has the same denominator as the first fraction?

OpenStudy (anonymous):

4, but then what would I do with the 3? It'd end u being \[\frac{ 3c-6d }{ 8c } + 3\]

OpenStudy (ddcamp):

For the three, multiply it by (8c/8c) to get \[\frac{ 3c-6d }{ 8c }-\frac{ 8c }{ 8c }3 = \frac{ 3c-6d }{ 8c }- \frac{ 24c }{ 8c }\]

OpenStudy (anonymous):

So the answer is \[\frac{ 21c-6d }{ 8c }\] ?

OpenStudy (anonymous):

then it would be 13c-6d?

OpenStudy (ddcamp):

Sorry, it's supposed to be +(24c/8c) The final answer should be (27c - 6d)/8c